A Bibliography of Publications of Barry Francis Smith

Barry Francis Smith

Tel: ?n/a? FAX: ?n/a?

E-mail: bsmith@mcs.anl.gov (Internet)

25 November 2011 Version 1.07

Abstract

This bibliography records publications of Barry Francis Smith.

Title word cross-reference

algorithm [DS89, Smi91, Smi92a, Smi93, WS90]. Algorithms [DSW94, DSW93]. Analysis [DSW94, DSW93]. analyze [DS89]. antisymmetries [DS89].

basis [WS90]. boundary [DS89].

case [DS89]. Coarse [CSZ94].

data [GS93]. data-structure-neutral [GS93]. Decomposition [CS93, GS92, Smi91, Smi92b, WS90]. design [GS93]. Dimensions [DSW94, GS92, DSW93, Smi91, Smi92a, Smi93]. Domain [CS93, GS92, Smi91, Smi92b, WS90].

elasticity [Smi92b]. Elliptic [DSW94, DS89, Smi91]. Experiences [GS92]. Extensible [GS94].

Grids [CS93, CSZ94].

hierarchical [WS90].

implementation [Smi93]. interaction [DHH^{+}88]. Iterative [DSW94, DSW93, GS93, Smi92a, Smi93].

Libraries [GS94, GS93]. linear [GS93].

machines [DHH^{+}88]. matching [CSZ94]. Meshes [CSZ94]. Methods [CSZ94, GS92]. Multigrid [CS93, DS89].

neutral [GS93]. Non [CSZ94]. Non-matching [CSZ94]. Numerical [GS94]. numerics [DHH^{+}88]. optimal [Smi92b]. Overlapping [CSZ94, GS92].

parallel [DS89, Smi93]. Portable [GS94]. preconditioner [Smi92b]. problem [DS89]. Problems [DSW94, DSW93, Smi91, Smi92a, Smi92b, Smi93].

Scalable [GS94]. Schwarz [CSZ94, DSW93, DSW94, GS92]. solution [GS93]. sparse [GS93]. structure [GS93]. Substructuring [DSW94, DSW93, Smi92a, Smi93]. symmetries [DS89]. systems [GS93].

Three [DSW94, GS92, DSW93, Smi91, Smi92a, Smi93].

Unstructured [CS93, CSZ94]. Using [DS89, CSZ94, WS90].

value [DS89].

References

Chan:1993:MDD

[CS93] Tony F. Chan and Barry F. Smith.

Multigrid and domain decomposition on unstructured grids. Technical Report 93-42, UCLA, Dept. of Mathematics, 1993. To appear in the Proceedings of 1993 International Conference on Domain Decomposition.

Chan:1994:OSM

[CSZ94] Tony F. Chan, Barry F. Smith, and Jun Zou. Overlapping Schwarz methods on unstructured meshes using non-matching coarse grids. Technical Report 94-8, UCLA, Dept. of Mathematics, 1994.

Douglas:1988:INM

[DHH^{+}88] Craig C. Douglas, Michael E. Henderson, Susumu Horiguchi, Willard L. Miranker, Barry F. Smith, and Andrew Winkler. The interaction of numerics and machines. Informations-technik, 30: 83–98, 1988.

Douglas:1989:USA

[DS89] Craig C. Douglas and Barry F. Smith. Using symmetries and antisymmetries to analyze a parallel multigrid algorithm: the elliptic boundary value problem case. SIAM Journal on Numerical Analysis, 26(6):1439–1461, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dryja:1993:SAI

[DSW93] Maksymilian Dryja, Barry F. Smith, and Olof B. Widlund. Schwarz analysis of iterative substructuring algorithms for problems in three dimensions. Technical Report MCS-P250-0791, Argonne National Laboratory, Argonne, IL, USA, 1993. Published in [DSW94].

Dryja:1994:SAI

[DSW94] Maksymilian Dryja, Barry F. Smith, and Olof B. Widlund. Schwarz analysis of iterative substructuring algorithms for elliptic problems in three dimensions. SIAM Journal on Numerical Analysis, 31(6):1662–1694, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gropp:1992:EDD

[GS92] William D. Gropp and Barry F.

Smith. Experiences with domain decomposition in three dimensions: Overlapping Schwarz methods. In Domain Decomposition Methods in Science and Engineering, pages 323–333. Amer. Math. Soc., Providence, RI, USA, 1992. ISBN 0-8218-5158-6.

Gropp:1993:DDS

[GS93] William D. Gropp and Barry F.

Smith. The design of datastructure-neutral libraries for the iterative solution of sparse linear systems. Technical Report MCS-P356-0393, Argonne National Laboratory, Argonne, IL, USA, 1993.

Gropp:1994:SEP

[GS94] William D. Gropp and Barry F.

Smith. Scalable, extensible, and portable numerical libraries. In Proceedings of Scalable Parallel Libraries Conference, pages 87–93. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Smith:1991:DDA

[Smi91] Barry F. Smith. A domain decomposition algorithm for elliptic problems in three dimensions. Num. Math, 60:219–234, 1991.

Smith:1992:ISA

[Smi92a] Barry F. Smith. An iterative substructuring algorithm for problems in three dimensions. In Tony F. Chan, David E. Keyes, Gérard A. Meurant, Jeffrey S. Scroggs, and Robert G. Voigt, editors, Fifth International Symposium on Domain Decomposition Methods for Partial Differential Equations, pages 91– 98 (of xiii + 623). SIAM, Philadelphia, PA, USA, 1992. ISBN 0-89871-288-2. LCCN QA402.2 .I57 1991.

Smith:1992:ODD

[Smi92b] Barry F. Smith. An optimal domain decomposition preconditioner for elasticity problems. SIAM J. Sci. Stat. Comput., 13: 364–378, 1992.

Smith:1993:PII

[Smi93] Barry F. Smith. A parallel implementation of an iterative substructuring algorithm for problems in three dimensions. SIAM J. Sci. Stat. Comput., 14:406–423, 1993.

Widlund:1990:DDA

[WS90] Olof B. Widlund and Barry F. Smith. A domain decomposition algorithm using a hierarchical basis. SIAM J. Sci. Stat. Comput., 11:1212–1220, 1990.